OpenStudy (anonymous):

x^2 + 40 x^2 - x + 190 What is x ?

OpenStudy (chaise):

Are the two equations equal to one another?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=150

OpenStudy (anonymous):

Explain?

OpenStudy (chaise):

Have you written your question right? Your other question is very similiar.

OpenStudy (anonymous):

Because its the same.

OpenStudy (anonymous):

you jst need to group all variable terms on one side and all the constand on the other side of the equation

OpenStudy (anonymous):

constant*

OpenStudy (anonymous):

\[\huge x^2+40=x^2-x+190\] \[\huge 40=-x+190\] \[\huge -190+40=-x\] \[\huge -150=-x\] \[\huge x=150\]

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

but isnt it 50?

OpenStudy (anonymous):

because dont you square root it?

OpenStudy (anonymous):

square root what?

OpenStudy (anonymous):

150?

OpenStudy (anonymous):

why would you get the square root of that?? :|

OpenStudy (anonymous):

-150 = -x

OpenStudy (anonymous):

youll just divide both by -1 to end up with that answer :)

OpenStudy (anonymous):

to get rid of the negative

OpenStudy (anonymous):

do you get it now ? :)

OpenStudy (anonymous):

its -150 though you forgot the 2x